Griggs Company produces a single product with a current selling price of $170. Variable costs are $130 per unit, and fixed costs per month average $6,240. Management is considering increasing the selling price to a proposed $190 per unit. Assume that the variable cost per unit of the product and monthly fixed expenses will not change as a result of the proposed increase in selling price.

Hint: Treat each situation (current and proposed price) as separate potential scenarios when evaluating each question.

At the current selling price of $170 per unit, closest to what dollar volume of sales per month is necessary for Griggs to generate a monthly profit of $12,000? (Round the intermediate percentage to one decimal place.)

Multiple Choice

  • $51,063

  • $24,162

  • $77,617

  • $58,838
Answer #1

Contribution margin ratio = Contribution margin/Sales

= (170-130)/170 = 23.5%

Sales needed = (Fixed costs+target profit)/Contribution margin ratio

= (6240+12000)/23.5%

= 77,617